### v1

# SPHEREx Pre-Perihelion Mapping of $\mathrm{H_{2}O}$, $\mathrm{CO_{2}}$, and $\mathrm{CO}$ in Interstellar Object 3I/ATLAS

Abstract
From 01- to 15-Aug-2025UT, the SPHEREx spacecraft observed interstellar object 3I/ATLAS. Using $R=40\mathrm{-}130$ spectrophotometry at $\lambda=0.7\mathrm{-}5\,\mathrm{\mu m}$, light curves, spectra, and imaging of 3I were obtained. From these, robust detections of water gas emission at $2.7\mathrm{-}2.8\,\mathrm{\mu m}$ and $\mathrm{CO}_{2}$ gas at $4.23\mathrm{-}4.27\,\mathrm{\mu m}$ plus tentative detections of $^{13}\mathrm{CO}_{2}$ and $\mathrm{CO}$ gas were found. A slightly extended $\mathrm{H_{2}O}$ coma was detected, and a huge $\mathrm{CO}_{2}$ atmosphere of extending out to at least $4.2\times 10^{5}\,\mathrm{km}$ was discovered. Gas production rates for $\mathrm{H_{2}O}$, $^{12}\mathrm{CO}_{2}$, $^{13}\mathrm{CO}_{2}$, and $\mathrm{CO}$ were $Q_{gas}=3.2\times 10^{26}\pm 20\,\mathrm{\%}$, $1.6\times 10^{27}\pm 10\,\mathrm{\%}$, $1.3\times 10^{25}\pm 25\,\mathrm{\%}$, and $1.0\times 10^{26}\pm 25\,\mathrm{\%}$, respectively. Co-addition of all $\lambda=1.0\mathrm{-}1.5\,\mathrm{\mu m}$ scattered light continuum images from produced a high SNR image consistent with an unresolved source. The scattered light lightcurve showed $\lesssim 15\,\mathrm{\%}$ variability over the observation period. The absolute brightness of 3I at $1.0\mathrm{-}1.5\,\mathrm{\mu m}$ is consistent with a $2.5\,\mathrm{km}$ radius nucleus surrounded by a 100 times brighter coma. The $1.5\mathrm{-}4.0\,\mathrm{\mu m}$ continuum structure shows a strong feature commensurate with water ice absorption seen in KBOs and distant comets. The observed cometary behavior of 3I, including its preponderance of $\mathrm{CO}_{2}$ emission, lack of $\mathrm{CO}$ output, small size, and predominance of large icy chunks of material in a flux-dominant coma is reminiscent of the behavior of short period comet 103P/Hartley 2, target of the NASA Deep Impact extended mission in 2010 and a “hyperactive comet” near the end of its outgassing lifetime. This correspondence places 3I closer to barely- or non-active 1I/Oumuamua than primitive, ice rich 2I/Borisov, suggesting that ISOs are often highly thermally processed before ejection into the ISM.
[…]
facilities: SPHEREx, IRTF, SOLOsoftware: SPHEREx-Sky-Simulator (Crill et al., 2025), spiceypy (Annex et al., 2020), SPICE (Acton et al., 2018; Acton, 1996), kete (Dahlen et al., 2025), skyloc [^11] https://github.com/ysBach/skyloc, SPHEREx-SSO, astropy(Astropy Collaboration et al., 2013, 2018, 2022), ccdproc (Craig et al., 2025), photutils (Bradley et al., 2025),
show]carey.lisse@jhuapl.edu
1 Introduction
SPHEREx (Spectro-Photometer for the History of the Universe, Epoch of Reionization and Ices Explorer) is a new NASA mission launched into low-Earth, Sun-sychronous orbit (altitude $\approx 675\,\mathrm{km}$) on 11-Mar-2025 UT (Doré et al., 2016, 2018; Bock et al., 2025). Conducting a 102-band near-infrared spectrophotometric survey of the entire sky over the course of 2 years, it was designed primarily to study (1) inflationary cosmology, (2) the history of galaxy formation, and (3) the abundance of astrobiologically important molecular ices in planet-forming regions. SPHEREx will observe and catalog measurements of everything along a line of sight on the sky—from local solar system objects (asteroids, comets, planets, KBOs) to main sequence stars, to distant quasars. One of these objects, 3I/ATLAS (3I, hereafter), was recently discovered on 01-Jul-2025 (Denneau et al., 2025) and verified within the next few days as only the 3rd ever macroscopic body detected hurtling through the solar system from the depths of interstellar space. Termed “ISOs”, bodies like 3I represent unique samples of exosystem material delivered close to Earth “for free”, allowing in-depth remote sensing analysis of their size, shape, composition, and temporal behavior during the few months when they are passing through the inner solar system.
Analysis of 3I’s orbit and apparent motion on the sky [JPL Horizons Ref] with respect to the SPHEREx sky survey region (Solar elongation $\sim 90\arcdeg$; Bryan et al. (2025)) showed that the object would be visited $\sim~70$ times from 08- to 15-Aug-2025 during the normal survey (Fig. 1), with some gaps in wavelength coverage, especially around $3\,\mathrm{\mu m}$, due to 3I’s rapid apparent sky motion. The project then proceeded to design, test, and implement a special observing mode, only 2 months after the start of the mission’s all-sky survey, to augment the total number of visits to 3I up to 160 flux measurements while completely characterizing the object from $0.75\mathrm{-}5.0\,\mathrm{\mu m}$.
The initial results from the first $\sim 70$ SPHEREx normal survey visits to 3I, as published in Lisse et al. (2025), resulted in the discovery of a giant $\mathrm{CO}_{2}$ coma surrounding the nucleus and the presence of large ($100\,\mathrm{\mu m}$) water-ice rich grains in the coma. Scattered light was detected from the nucleus and coma grains at $\lambda\lesssim 2.3\,\mathrm{\mu m}$, but only 3-$\sigma$ upper limits were placed on the emission rates of $\mathrm{H_{2}O}$ and $\mathrm{CO}$ gas.
In this paper, with more than double the total measurements and a much better data reduction and data calibration, we extend the preliminary SPHEREx discoveries to include a robust detection of the water gas emission feature at $2.7\mathrm{-}2.8\,\mathrm{\mu m}$ and a marginal detection of $\mathrm{CO}$ gas emission at $4.7\,\mathrm{\mu m}$.
In addition, the detected $^{12}\mathrm{CO}_{2}$ line emission at $4.25\mathrm{-}4.27\,\mathrm{\mu m}$ was so strong that we searched for substructure in it. While we were not able to resolve the $4.25$ vs $4.27\,\mathrm{\mu m}$ doublet at the peak, we were able to spectrally distinguish a shoulder at $4.37\,\mathrm{\mu m}$ where emission due to $^{13}\mathrm{CO}_{2}$ is found. Assuming the same distribution pattern as for the $^{12}\mathrm{CO}_{2}$ and thus the same aperture size, we find the abundance ratio of $^{13}\mathrm{CO}_{2}$ to $^{12}\mathrm{CO}_{2}$ to be the same, within the error bars of the measurement, as reported by Cordiner et al. (2025) using JWST NIRSPEC IFU spectroscopy.
SPHEREx’s unique observational strengths lie in its ability to spectrophotometrically image an object at $0.75\mathrm{-}5.0\,\mathrm{\mu m}$ over many weeks and on large angular scales. SPHEREx’s spectrophotometric quality is attested to by its overall excellent 3I spectral match to the published $0.7\mathrm{-}5.3\,\mathrm{\mu m}$ JWST spectrum (Cordiner et al., 2025); where they differ, in the gas emission lines, it is always SPHEREx’s measurement that is the larger due to its much larger projected observing aperture (Fig. 2). (It is important to note that the area of the SPHEREx central pixel is 4 times the area of the entire JWST NIRSPEC IFU field of view.) SPHEREx’s temporal capabilities are demonstrated by the the 12 day long dust continuum reflectance lightcurve (Fig. 1c) and its large field image quality by $\sim 6\arcmin$ wide $\mathrm{CO}_{2}$ coma detection shown in Fig. 3.
The chief systematic found for the SPHEREx 3I observations is confusion with background stars. Other effects, like motion with respect to the Sun and Earth, and changes in phase angle, were small and easily corrected for. Object variability, which can introduce large uncertainties in SPHEREx spectrophotometry obtained piecemeal over 15 days, was not an issue due to the lack of any observed lightcurve variation greater than 15% found (Fig. 1c) and in reports by other groups (Kareta et al., 2025; Chandler et al., 2025). This lack of variation, in turn, is likely due to the dominance of the observed flux by the slowly time-varying smooth coma, which can mask any rapid projected surface area variations of an asymmetric rotating nucleus.
In this Letter we describe the complete set of August 2025 SPHEREx observations of 3I, and their data quality, reduction, and analysis. Because of these observations, we now understand that the extended optical morphology seen for 3I since its discovery is the product of robust $\mathrm{CO}_{2}$ gas sublimation that was undetectable in the first UVIS characterizations of the object reported in July 2025. We also find a very $\mathrm{CO}$-poor but $\mathrm{H_{2}O}$-ice rich cometary object with its water gas production just beginning to turn on, and evidence for thermal processing of the body as seen in highly thermalized inner solar system comets. We find good concordance with other recent imaging and spectroscopic measures of the object. We use these findings to argue that 3I must have been well-thermally processed by many close passages of its home star before ejection by a co-orbiting Neptune to stellar mass body, and that such thermal processing happens more rapidly than exosystems dynamically relax.
2 Observations
SPHEREx obtained 160 frames with 3I within the field of view from 2025-08-01 UT 08:01:07 to 2025-08-15 UT 12:43:48 (midpoints of exposures), including 94 exposures produced by deliberately adjusting the survey planning algorithm to place 3I in the FoV. A subset of these were downlinked and processed promptly, leading to our express publication in Lisse et al. (2025). In this work, we analyze all 160 frames (Table 1). The orbital geometry of the observation and the sky mosaic images are shown in Fig. 1.
Figure 1: Motion of 3I. (a) Color composite from SPHEREx data at 1.185, 1.716, and 2.194 $\mathrm{\mu m}$ for red, green, and blue, respectively. The trajectory of 3I is overplotted, with observed positions marked by crosses. The starting and ending points of our dataset are indicated separately (see legend). (b) Orbital motion of 3I (black) projected onto the ecliptic plane. The orbits of Earth, Mars, and Jupiter are shown in blue, red, and purple, respectively. Numbers indicate the positions of the bodies at the first day of each month. 3I’s retrograde motion and perihelion are marked with arrows. (c) Lightcurve of 3I produced by taking all $\lambda=1.0\mathrm{-}1.8\,\mathrm{\mu m}$ SPHEREx flux measurements and dividing them by the JWST flux (Fig. 2b; Cordiner et al. 2025) to leave only lightcurve variations. Also included are SOLO clear-filter brightness measurements scaled to match their median to that of the SPHEREx/JWST ratio. The resulting lightcurve amplitude is within $\pm 15\,\mathrm{\%}$, so the derived reflectance should be correct to within this range considering any lightcurve modulation effects.
Ancillary lightcurve observations were obtained using the remote/robotic Solar system Objects Light-curve Observatory (SOLO; B. Lim et al. in prep), a Celestron LLC’s RASA11 telescope [^1] An 11-inch reflecting telescope manufactured by Celestron LLC located at Sierra Remote Observatory [^2] IAU observatory code G80, $37\arcdeg 4\arcmin 13.39\arcsec$ N, $119\arcdeg 24\arcmin 45.72\arcsec$ W, CA, USA in order to independently check the SPHEREx time domain trending. Combined with the currently installed imaging sensor at the prime focus, KL4040333Manufactured by FLI Kepler, a 4k by 4k CMOS Back Illuminated (BI) camera, and a UV-cut clear filter, the total throughput of the SOLO system as of writing is effectively sensitive to $\sim 0.4\mathrm{-}1.0\,\mathrm{\mu m}$, with a response closely comparable to the Gaia G band. The results are shown in Fig. 1c, and are consistent with the low (if any) level of variability detected by SPHEREx.
Ancillary spectral observations were used to independently check the SPHEREx spectral measurements were obtained at the NASA Infrared Telescope Facility (IRTF) on the nights of 09 and 10-Aug-2025 using the SpeX instrument in $R\sim 50$, $\lambda=0.7\mathrm{-}2.5\,\mathrm{\mu m}$ prism spectroscopy mode (Fig. 2). These measurements were vital for verifying initial SPHEREx spectrophotometric data reductions and quickly bootstrapping early photometric measures, and the two sets of mesurements were found to be closely consistent, especially the strong dropoff past $2.3\,\mathrm{\mu m}$ in the IRTF spectrum that was initially believed to be an instrumental artifact near the end of the spectral order.
3 Data Reduction
We first simulated the exposures using the SPHEREx reference catalog (Y. Yang et al., in prep.) and the SPHEREx Sky Simulator (Crill et al., 2025). The simulated images, which do not have 3I in them, were then subtracted from the calibrated raw Level 2b (l2b) image data provided by the SPHEREx science pipeline in Quick Release 2 (QR2; SPHEREx Explanatory Supplement). This was extremely helpful in removing the estimated flux of field objects within 3I’s aperture, flagging less reliable images and extendedness artifacts, and significantly improved the final spectrum.
For signal extraction, we applied simple circular aperture photometry with a $12\arcsec$ radius ($\sim 2$ pixels, depending on the location in the FoV due to the distortion). The residual local sky background was estimated from a circular annulus with inner and outer radii of $\approx 1\arcmin$ and $3\arcmin$, respectively. The photometry at the same celestial location from the original l2b image, simulated image, and the simulation-subtracted image are $F_{\mathrm{2b}}$, $F_{\mathrm{sim}}$, and $F_{\lambda}=F_{\mathrm{2b}}-F_{\mathrm{sim}}$, respectively.
We verified that the results are robust against variations in the annulus size and different sky estimation algorithms. As an additional null test, we measured the single pixel value minus the estimated sky from the surrounding annulus for several pixels that are neither flagged as bad nor expected to contain source signals (based on the reference catalog) within each image. These values are consistent with zero, with scatter comparable to the expected pixel variance. This confirms that the sky estimation is reliable at the pixel-variance level.
The uncertainty of $F_{\lambda}$, $dF_{\lambda}$, is obtained based on the aperture sum of the associated variance map of the l2b image, quadratically combined with the local sky-estimation error. The variance map includes both readout and Poisson noise terms (Crill et al., 2025; Akeson et al., 2025). Since this does not take into account the uncertainty of the fluxes of field objects, it should be regarded as a lower limit; however, it should be close to the true value when sidereal object contamination within the aperture is small.
We define the contamination factor as
$F_{\mathrm{field}}\equiv F_{\mathrm{sim}}/F_{2b}~.$ (1)
For example, if a contaminant contributes half the flux of 3I, then we expect $F_{\mathrm{field}}=0.5/(1+0.5)\approx 0.3$. The higher this number, the larger the measurement uncertainty.
Among the 160 exposures, 9 were flagged as “a” and are completely discarded because at least one unrecoverable pixel (cold, hot, non-functional, early transient, and early overflow) is within the aperture, and it is unlikely that they can ever be improved. Flag “b” is given to 59, and they are regarded as less reliable as of writing because of strong background source contamination expected and/or calibration ($F_{\mathrm{field}}>30$ %, object is on the edge of detectors, etc). Finally, 11 were flagged “c”, which means we see a non-negligible contamination likely from background objects identified by visual inspection, that is, the flux interpolated from the reference catalog does not fully represent the sidereal sources near 3I. In principle, fluxes from exposures with flags “b” and “c” could be improved once SPHEREx obtains images of the same sky regions at the same wavelengths, when suitable “reference” images will be available for subtraction.
4 Results
The direct observables produced by SPHEREx spectrophotometric imaging of 3I in early August 2025 fall into 3 broad categories: temporal lightcurves, spectrophotometry, and imaging.
4.1 Lightcurves
Lightcurve analysis of 3I’s temporal behavior (Fig. 1) shows very little variability of $\lesssim 15\,\mathrm{\%}$ from hours to days timescales, which is also confirmed by independent SOLO observations. This upper limit of variation is also consistent with previous reports (Seligman et al., 2025; Kareta et al., 2025; Beniyama, 2025). The fortuitous lack of temporal variability, explained by an object with a small, likely dark nucleus surrounded by a very bright and high surface area coma (see Section 4.6 below), meant that we could quickly cull out any background contaminated images and proceed straight to analysis of the data.
4.2 Spectra
The measured fluxes of 3I are scaled to the geometry of $r_{\mathrm{obs}}=r_{\mathrm{hel}}=1\,\mathrm{au}$ assuming $F_{\lambda}\propto r_{\mathrm{hel}}^{-2}r_{\mathrm{obs}}^{-2}$ (Fig. 2a). The reflectance spectrum is derived by dividing the SPHEREx measurements by the solar flux. We can break up the spectrum into 5 main regions: the $0.7\mathrm{-}1.5\,\mathrm{\mu m}$ scattered light region; the $1.5\mathrm{-}2.3\,\mathrm{\mu m}$ mixed organics plus water absorption region; the $2.4\mathrm{-}3.3\,\mathrm{\mu m}$ water ice absorption region; the $3.2\mathrm{-}3.7\,\mathrm{\mu m}$ organics absorption region; and the $3.7\mathrm{-}5.0\,\mathrm{\mu m}$ mixed region. Superimposed on top of the continua are weak gas emission features due to $\mathrm{H_{2}O}$ at $2.7\mathrm{-}2.8\,\mathrm{\mu m}$ and $\mathrm{CO}$ at $4.6\mathrm{-}4.8\,\mathrm{\mu m}$, and a strong $\mathrm{CO}_{2}$ emission feature at $4.2\mathrm{-}4.4\,\mathrm{\mu m}$. The $\mathrm{CO}_{2}$ feature is so strong, in fact, that we believe we have detected not only normal $^{12}\mathrm{CO}_{2}$ centered in a doublet at $4.25$ and $4.27\,\mathrm{\mu m}$, but also a small side shoulder due to $^{13}\mathrm{CO}_{2}$ centered at $4.31\,\mathrm{\mu m}$. The agreement in spectral shape with the JWST spectrum of Cordiner et al. (2025) taken in one visit during the middle of our observational window is excellent; the main disagreement between the two spectral datasets is in the amplitude of the gas emission lines, with SPHEREx’s consistently being higher due to the much larger effective aperture of its measurements compared to the $3\arcsec\times 3\arcsec$ FOV of the JWST NIRSPEC IFU (approximately 1/4 the area of one SPHEREx pixel).
Figure 2: (a) The flux density of the measurements, scaled to the geometry of $r_{\mathrm{obs}}=r_{\mathrm{hel}}=1\,\mathrm{au}$. The JWST measurements (Cordiner et al., 2025) are scaled by a constant factor of 1.5 to match the SPHEREx measurements. For flags, see Table 1 and text for explanation. The skyblue points are the JWST measurements, convolved with the pixel spectral throughput at the SPHEREx pixel for that observation, using the SPHEREx Sky Simulator (Hui et al., 2024; Crill et al., 2025). Blue shades indicate important spectral features. Colors of “good” measurements indicate $F_{\mathrm{field}}$ (Eq. 1). (b) The distance-corrected ratio of the SPHEREx to JWST flux, showing that SPHEREx measurements are systematically higher than JWST. This indicates the spatial extendedness of the object, 3I. The error-bars are not shown for $S/N<2$ data. (c) The reflectance, normalized at $1.2\,\mathrm{\mu m}$ for SPHEREx, JWST, and the IRTF measurements. The blue line shows the moving median of the IRTF data with a $0.2\,\mathrm{\mu m}$ window.
Because the $\mathrm{CO}_{2}$ coma is definitely extended beyond $1\arcmin$ (Fig. 3), the nominal aperture photometry will give only a lower limit of the flux. Integrating the total flux using the radial profile is discussed below.
Figure 3: Median combined images ($\approx 5\arcmin\times 5\arcmin$) and radial profiles. Left column: The stacked images, after simulation and background subtractions, in MJy/sr unit (see scale bars). The orange bars at the bottom indicate aperture diameter ($24\arcsec$) and $1\arcmin$. Middle column: The radial profile of 3I in the stacked image (green; blue if the signal is above the 1-$\sigma$ scatter), mean (gray) and median (black) of 30 brightest field stars within $\lesssim 10\arcmin$, respectively. Each profile is normalized to the central pixel. The red solid and dashed lines are for $\rho^{-1.5}$ and $\rho^{-1}$ profiles, respectively, for the radius $\rho$. The nominal aperture radius (orange vertical lines) and corresponding length at 3I (title) are shown. Right column: Same as the left column but convolved with a standard deviation of 1.5-pixel gaussian filter for visualization. In the left and right columns, celestial North (white thick arrow), East (white thin arrow), anti-solar (red solid) and anti-velocity (yellow dashed) vectors are shown in the top left corner. The white and red contours (also shown in the color bar to the right of each panel) are 2- and 5-$\sigma$ above the background, respectively.
4.3 Gas Production Rates
We adopt a production rate $Q_{\mathrm{gas}}$ via
$Q_{\mathrm{gas}}=2v_{\mathrm{gas}}rN_{\mathrm{gas}}$ (2)
[…]
$N_{\mathrm{gas}}=\frac{4r_{\mathrm{obs}}^{2}}{r^{2}}\frac{F_{\nu}\Delta\lambda}{gh\lambda}\left(\frac{r_{\mathrm{hel}}}{1\,\mathrm{au}}\right)^{2}$ (3)
where $F_{\nu}$ is the average flux density (i.e. power per area per frequency bandwidth) in the band from the gas; $r_{\mathrm{hel}}=3.2\,\mathrm{au}$ and $r_{\mathrm{obs}}=2.6\,\mathrm{au}$ are the heliocentric and observer-centric distances to 3I, respectively; $g$ is the fluorescence efficiency at 1 au for the emission band in question, and we use $g=2.9\times 10^{-4}$, $2.6\times 10^{-3}$, and $2.6\times 10^{-4}$ photons per second per molecule for $\mathrm{H_{2}O}$ $2.85\,\mathrm{\mu m}$, $\mathrm{CO}_{2}$ $4.25\,\mathrm{\mu m}$, and $\mathrm{CO}$ $4.7\,\mathrm{\mu m}$ emission bands, respectively (Crovisier et al., 1997); $h$ is Planck’s constant; $\Delta\lambda$ is the bandwidth for the narrow-band assumption, and we use $0.08\,\mathrm{\mu m}$ for $\mathrm{H_{2}O}$ and $0.037\,\mathrm{\mu m}$ for all others; and $\lambda$ is the wavelength of the band.
Given the SPHEREx $\mathrm{H_{2}O}$, $^{12}\mathrm{CO}_{2}$, $^{13}\mathrm{CO}_{2}$, and CO gas emission fluxes of 58, 8610, 70, and 59 mJy measured in 2 pixel radius effective apertures and corrected for 3I’s relative position to rh = 1 au, Delta = 1 au (Fig. 2), we derive gas production rates of $Q_{\mathrm{\mathrm{H_{2}O}}}=3.2\times 10^{26}\,\mathrm{molec/sec}\pm 20\,\mathrm{\%}$, $Q_{\mathrm{{}^{12}\mathrm{CO}_{2}}}=1.6\times 10^{27}\,\mathrm{molec/sec}\pm 10\%$, $Q_{\mathrm{{}^{13}\mathrm{CO}_{2}}}=1.3\times 10^{25}\,\mathrm{molec/sec}\pm 25\%$, and $Q_{\mathrm{\mathrm{CO}}}=1.0\times 10^{26}\,\mathrm{molec/sec}\pm 25\,\mathrm{\%}$. The rates of $\mathrm{CO}_{2}$ and $\mathrm{H_{2}O}$ emission are consistent with the activity of inbound thermally processed short-period Solar System comets at $3.2\,\mathrm{au}$ (Harrington Pinto et al., 2022), where water gas emission is “not fully on yet”, and the ratio of $Q_{\mathrm{{}^{13}\mathrm{CO}_{2}}}/Q_{\mathrm{{}^{12}\mathrm{CO}_{2}}}\sim 1/100$ is consistent with the ISM value 0f 70, although the $\mathrm{CO}$ production rate at $0.1Q_{\mathrm{CO}_{2}}$ is very low (see Discussion section). These updated Qgas values are also consistent with the preliminary values reported for the first 70 visits by Lisse et al. (2025), and by Cordiner et al. (2025) from their JWST observations of 06-Aug-2025 UT taken roughly halfway through the SPHEREx observing run.
4.4 Images
We extracted cutout images at the specified wavelengths, with the few pixels flagged as cold, hot, non-functional, or otherwise erroneous masked. The masked images were then median-combined, and the background removed by subtracting the smoothed 10 by 50 pixels rectangular regions, centered at 75 pixels away to the left and right of the center of 3I. The resulting stacked, background removed SPHEREx near-infrared images of 3I are shown in Fig. 3.
Images at different wavelength ranges probe the distribution of dust and important spectral features by molecules (Sect. 4.2). The large extension seen in $^{12}\mathrm{CO}_{2}$ versus $\mathrm{H_{2}O}$ is not surprising given the orders of magnitude higher measured flux emission from $\mathrm{CO}_{2}$ gas, the object’s location well within the $\sim 12\,\mathrm{au}$ $\mathrm{CO}_{2}$ ice line of the solar system, and the long lifetime ($\sim 6$ days) vs photolysis and SWCX compared to water ($\sim 1$ day). However, the falloff of the $\mathrm{CO}_{2}$ coma with respect to projected radial distance $\rho$ is definitely faster/steeper than $1/\rho$, implying that there is either acceleration of the outflow occurring or there is destruction of the outflowing $\mathrm{CO}_{2}$ going on.
The lack of a bright water gas coma is puzzling as 3I was not far outside the Solar system’s “water ice line” at $r_{\mathrm{hel}}=2.5\,\mathrm{au}$ during the observations. We are led to conclude that 3I was emitting large chunks of mixed $\mathrm{CO}_{2}$$\mathrm{H_{2}O}$ ice into its coma at the time of observation, and that evaporative cooling of $\mathrm{CO}_{2}$ was pinning the chunks’ temperature at $\sim 120\,\mathrm{K}$ and greatly suppressing the $\mathrm{H_{2}O}$ ice’s vapor pressure (Lisse et al., 2021). A corollary of this argument is that once 3I comes well within the water ice line of the solar system at $2.5\,\mathrm{au}$, its outgassing behavior will markedly change and increase as majority matrix water ice itself starts evaporating rapidly. This will destroy many of the large chunks of ice in the coma, changing it over to a small particle dominated cloud with a more anti-solar pointing tail, while driving the sublimation front inward towards the nucleus.
4.5 Radial profiles
[…]
Only at wavelengths characteristic of emission from $\mathrm{H_{2}O}$ and $^{12}\mathrm{CO}_{2}$ do we see significant extension above stellar; the profile for $\mathrm{H_{2}O}$ is noisy enough that we can’t say much more about its power-law slope. By contrast, the $^{12}\mathrm{CO}_{2}$ profile slope is very well defined, and clearly falls off more quickly than a $\rho^{-1}$ coma, more like a $\rho^{-3/2}$ coma. This is indicative of either significant photolytic + charge exchange destruction of the $\mathrm{CO}_{2}$, or significant acceleration of the $\mathrm{CO}_{2}$. Given that the $\mathrm{CO}_{2}$ coma extends out to $\sim 3\arcmin=384,000\,\mathrm{km}$ at $3.2\,\mathrm{au}$, and $\mathrm{CO}_{2}$ gas molecules have a photolytic lifetime of $\sim 6\,\mathrm{days}=526,000\,\mathrm{s}$ and move at average speed of $0.8\,\mathrm{km/sec}$, it is hard to see how many of them could be getting destroyed close to the nucleus. Instead they are likely accelerating due to significant pressure in the gas releasing coma, an explanation supported by noting that the $\mathrm{CO}_{2}$ coma gas profiles of Fig. 3 also appears quite flat for the first $\sim 15\arcsec$ ($32,000\,\mathrm{km}$) or so before bending past a sublimation front to follow a $\rho^{-3/2}$ trend at larger distances.
4.6 Nucleus Size
Since the scattered light signal from 3I is unresolved spatially versus a point source by SPHEREx, we analyze its importance by taking limiting cases. Firstly, if we assume all the flux were to originate as sunlight scattered off of a nuclear surface of geometric albedo $p_{\mathrm{V}}=0.04$ at the phase angles $\alpha=16.3\mathrm{-}17.2\arcdeg$ of 3I during the observation, we find an effective spherical radius of the object of $\sim 23\,\mathrm{km}$ (Lisse et al., 2025). However, we know this assumption is patently false, as Hubble Space Telescope observations of the object have placed a firm upper limit, after fitting and tracing the coma flux back to the central PSF pixels, of no more than an $2.5\,\mathrm{km}$ radius, $p_{\mathrm{V}}=0.04$ point source residual. Further, no significant lightcurve has ever been measured for 3I, and a large asymmetric rotating nucleus, as all comets observed to date have shown, will produce significant lightcurve variations. And finally, the huge $\mathrm{CO}_{2}$ coma mapped by SPHEREx shows a flat-topped structure within 18” (3 pixels or 42,000 km) of the nucleus. The solution to all these issues is the same: the presence of a small, rotating, dark nucleus surrounded by a stable, approximately 100 times brighter coma. (Cometary comae are typically static in brightness over hours to days unless the nucleus produces a huge outburst of emission or has a bright focused region of jet-like activity on its surface). Solar system comets where the nucleus is greatly outshown by its coma are rare but not unheard of; they are the handful of small, “hyperactive” comets like 41P/Wirtanen and 103P/Hartley 2 (103P, hereafter) with so much icy dust in their comae that the coma solid material has more surface area than the nucleus and out-produces the nucleus in gas emission. These comets thus require high amounts of large, long lived icy particle emission, and further appear to be $\mathrm{CO}_{2}$ gas dominated and near the end of their actively outgassing lives ((Lisse et al., 2009); see the Discussion Section for further arguments for the connection between 3I and 103P). This was just the kind of behavior 3I was exhibiting during the early August 2025 SPHEREx observations reported here.
5 Discussion
All of the observations reported here (the giant $\mathrm{CO}_{2}$ coma, the lack of any observed lightcurve variability, the prevalence of large coma grains, and the 100x greater coma scattered light flux than a $2.5\,\mathrm{km}$ radius, 0.04 geometric albedo spherical object would produce) are consistent with an object spectrally dominated by bright, $\mathrm{CO}_{2}$- and water ice-rich coma material. The low amount of $\mathrm{H_{2}O}$ gas production observed by SPHEREx in August 2025 can be explained by the object’s having only recently arrived from the cold depths of interstellar space and its remoteness from the Sun (outside the water ice line between $3.4$ and $3.1\,\mathrm{au}$) during the observational period (Lisse et al., 2021, 2022a). The lack of strong $\mathrm{CO}$ emission is slightly surprising, as the $\mathrm{CO}$-ice line in the solar system lies at $\sim 40\,\mathrm{au}$ from the Sun, until one realizes it is an indication of hypervolatile depletion.
A strong predominance of $\mathrm{CO}_{2}$ gas emission coupled with $\mathrm{CO}$ hypervolatile depletion has been seen in solar system hyperactive comets like 103P (A’Hearn et al., 2012, $Q_{\mathrm{CO}_{2}}/Q_{\mathrm{H_{2}O}}\sim 20\,\mathrm{\%}$; $Q_{\mathrm{CO}}/Q_{\mathrm{H_{2}O}}\sim 0.15\,\mathrm{\%}$) and 46P/Wirtanen (McKay et al., 2021, $Q_{\mathrm{CO}_{2}}/Q_{\mathrm{H_{2}O}}\sim 15\,\mathrm{\%}$; $Q_{\mathrm{CO}}/Q_{\mathrm{H_{2}O}}<0.5\,\mathrm{\%}$). In fact, 103P is the comet with the lowest known ratio of $\mathrm{CO}$ to $\mathrm{CO}_{2}$ (Harrington Pinto et al., 2022, see Fig. 9a). Hyperactive comets are defined as comets with gas production rates many times what sublimation from their nucleus surfaces can/should be able to support; their gas production is instead produced primarily from large, long-lived icy grains emitted into the surrounding coma by the nucleus.
Figure 4: (a) $\mathrm{CO}/\mathrm{CO}_{2}$ gas production rate ratios of solar system comets (Harrington Pinto et al., 2022). The $\mathrm{CO}$ or $\mathrm{CO}_{2}$ dominance criteria of $1\pm 0.33$ is indicated as horizontal shadow (Harrington Pinto et al., 2022). The ratio of 3I found by SPHERExis $\lesssim 0.013$, making it extremely $\mathrm{CO}$ poor and equivalent to the most thermally processed comet in the 2022 survey, 103P. (b) Pre-perihelion optical lightcurves observed for comet 103P during the 2010–2011 apparition and 3Iin May - Oct 2025. Published data from (Jewitt & Luu, 2025; Tonry et al., 2025; Ye et al., 2025) have been shifted to match the absolute magnitude of Minor Planet Center (MPC) Database [^7] https://www.minorplanetcenter.net/db_search/show_object?object_id=3I for 3I and https://www.minorplanetcenter.net/db_search/show_object?object_id=103P for 103P. , the T-mag from JPL Horizons ephemerides service [^8] https://ssd.jpl.nasa.gov/horizons/app.html, and the data kindly shared by S. Yoshida (2025999http://www.aerith.net/comet/catalog/0003I/2025N1.html, priv. comm.). There is very good agreement between the two inbound lightcurve trends. An inherent spread of $\sim 0.5\,\mathrm{mag}$ in the measurements between nights and between observers can be seen. (c) Close flyby imaging of the $\sim 1.5\times 0.6\times 0.6\,\mathrm{km}$ comet 103P nucleus as obtained by the Deep Impact Extended mission (A’Hearn et al., 2011). A marked bilobate nuclear morphology was seen, with unusually large ($m$ to dm) chunks of $\mathrm{CO}_{2}$ rich ice and gas emanating from the end of the smaller lobe and the middle of the larger lobe. The surrounding coma, filled with large long lived icy grains, was found to be at least 5 times brighter and more actively emitting gas than the nucleus. While the nucleus of 3I has never been resolved, given its similar behavior, it could possibly share similar morphological characteristics. (d) Schematic of the 3I object derived from SPHEREx $\mathrm{CO}_{2}$ photometric mapping. The $2.5\,\mathrm{km}$ radius nucleus (Jewitt et al., 2025) was not resolved at the SPHEREx pixel scale of $6.15\arcsec\approx 11,600\,\mathrm{km}$ at $2.6\,\mathrm{au}$ distance, but a notional nucleus image has been included at the origin to give the reader its location and a sense of scale. Two distances have been marked off: The $\sim 38,000\,\mathrm{km}$ (white dashed circle) inside of which we think coma ice is actively sourcing $\mathrm{CO}_{2}$ by sublimation; and $\sim 340,000\,\mathrm{km}$ (orange dashed oval), the minimum radial distance to which we detect $\mathrm{CO}_{2}$ gas at 2-$\sigma$ level.
The connection between 3I and 103P is bolstered by reports of 3I’s abundant C2 and CN emission, putting it akin to the spectral Class A comets of Fink 2009’s spectral survey, which also contains hyperactive comets 103P and 46P/Wirtanen. Further, comparison of the inbound visible lightcurve vs heliocentric distance for the 103P 2010 apparition to the current inbound lightcurve vs distance for 3I (Fig. 9b) shows an excellent match (Figure 4a).
103P was especially well studied during the close ($\sim 1000\,\mathrm{km}$ minimum distance) flyby of the NASA Deep Impact Extended mission (A’Hearn et al., 2011, in November 2010), which imaged large chunks of $\mathrm{CO}_{2}$ rich material being ejected “in a reverse snow” from one of the comet’s lobes and measured a coma $\sim 5\times$ brighter than the nucleus in total scattered light flux (Figure 4c). Because of 103P’s small size but near typical gas production rate, it has been argued to be near the end of its life as a comet in the inner solar system (Lisse et al., 2009). The large chunks of $\mathrm{CO}_{2}$-rich material are explained as primordial remnants of solid $\mathrm{CO}_{2}$ ice phases protected down deep near the center of an originally much larger nuclear body which has lost its much more hypervolatile ice phases (e.g., $\mathrm{CO}$, $\mathrm{N_{2}}$, and $\mathrm{CH_{4}}$). The imaging results from the Deep Impact extended mission suggest that the regions of its nucleus emitting $\mathrm{CO}_{2}$-rich ice are separated from regions producing water gas, and the source ices may be physically separated.
5.1 Other Hypotheses
There are several other hypotheses that have been proposed to explain 3I’s physical behavior. For example, it has been proposed that 3I originated in a very $\mathrm{CO}$-poor but $\mathrm{CO}_{2}$-rich exosystem (i.e., a/the nature hypothesis to our evolutionary nurture proposal of advanced thermal evolution). We find this explanation problematic, as this would imply a system with a host star C/O ratio <1/10<1/10 the Sun’s, and a dearth of reduced aliphatic and aromatic species as well. [C/Fe] and [C/H] <1.0<-1.0 for a nearby Milky Way star is highly improbable (Hinkel et al., 2017; Lisse et al., 2017; Mészáros et al., 2025)], and the aliphatic and PAH absorption features for 3I in the JWST NIRSPEC spectrum of Cordiner et al. (2025) look to be of reasonably normal strength.
It has also been proposed that 3I was sourced from an exosystem with near-solar composition, and it began as the same; but that galactic cosmic-rays preferentially transformed its near-surface $\mathrm{CO}$ ices into $\mathrm{CO}_{2}$ species via radiolytic driven production of OH, H, and O and subsequent oxidation of all $\mathrm{CO}$ to $\mathrm{CO}_{2}$. This proposed process would also tend to make $\mathrm{H_{2}CO}$, $\mathrm{CH_{3}OH}$, $\mathrm{HCOOH}$, as well, which may be consistent with current observations of 3I. However it is not consistent with the work of Moore et al. (2003), which showed that particle irradiation of either pure $\mathrm{CO}$ or $\mathrm{CO}_{2}$ in water ice tends to produce a mixture of the two species with abundance ratio on the order of unity, nor is it consistent with the $\mathrm{CO}/\mathrm{CO}_{2}$ ratios found in solar system Oort cloud comets, which spend $99\,\mathrm{\%}$ of their lifetimes in the ISM galactic cosmic ray radiation field, and emit abundant $\mathrm{CO}$. Close passages of Oort cloud comets by a star (the Sun), which 3I may never have experienced once its surface cr alteration began, would serve to drive off $\mathrm{CO}$ preferentially vs $\mathrm{CO}_{2}$, yet the $\mathrm{CO}/\mathrm{CO}_{2}$ abundance ratio remains high in these objects (Harrington Pinto et al., 2022).
We know that 3I is undergoing thermally driven evolution, as evidenced by the rate at which it is emitting dust and gas is trending like $r_{\mathrm{hel}}^{-3.8}$ after correcting for any observer-3I geometry (Jewitt & Luu, 2025), faster than the $r_{\mathrm{hel}}^{-2}$ trend for energy received from the sun is changing. We also know it is thermally evolving as we have watched its dust coma plus tail morphology go from large particle, sunward flying grains at discovery outside the water ice line, to the current small particle dominated, antisolar dust grains, as 3I’s the water production rate has picked up markedly inside the water ice line. We therefore consider a thermal evolutionary hypothesis for its behavior as not only highly likely, but necessary to consider.
5.2 Home System Implications
A heavily thermally processed state for 3I has a number of important implications. Firstly, by analogy to 103P’s behavior, it tells us that 3I has undergone 100s to 1,000s of close passages by its natal host/birth star (we reject the possibility that 3I has undergone 100’s to 1000’s of close passages of random galactic stars on its passsge to the solar system; this rate of close passage for a single object, even one as old as the galaxy, would imply a huge rate of interstellar object passages & detections in our solar system, which is not seen). These close thermally altering passages could have been accomplished either as a primordial body formed in-place near its birth system’s $\mathrm{CO}_{2}$ ice line (to ensure minimal CO but extensive labile $\mathrm{CO}_{2}$ ice incorporation) and quickly ejected soon after PPD disk clearing, or more classically, as in our own mature system, as an ice-rich KBO that became scattered into its inner system for a few hundred near-star orbits before ejection.
It also implies that 3I had a close encounter with a massive body that ejected it from its home system. Very simply, the maximum $\Delta v$ (velocity change) a close planetary flyby can impart is approximately the escape velocity for the planet. In our solar system, with $V_{\mathrm{esc}}\sim 25\,\mathrm{km/s}$ at $2.5\,\mathrm{au}$ from the Sun, only Jupiter ($V_{\mathrm{esc}}\sim 65\,\mathrm{km/sec}$), Saturn ($35.5\,\mathrm{km/sec}$), and barely Neptune ($25.5\,\mathrm{km/sec}$) could eject objects directly. Allowing for the possibility of a smaller mass birth star than the Sun definitely means that an inner planet as small as Uranus or Neptune could serve as an ejector, but this is about the minimum mass required because the water ice line retreats rapidly inward as the host star mass decreases ($V_{\mathrm{esc}}$ from the ice line to the ISM is actually larger for an M8V dwarfs like TRAPPIST-1, $\sim 34\,\mathrm{km/sec}$).
Objects commonly in close orbit around the host stars that are more massive than planets, like brown dwarfs, main sequence stars, AGB stars, white dwarfs, neutron stars, and black holes have a large enough $V_{\mathrm{esc}}$ to cause ejection from the ice line of a system, so rather than an inner massive planet, 3I could have come from a close-in binary system (Jackson et al. 2018, 2018MNRAS.478L..49J).
5.3 ISO Evolution
Finally we end with one last piece of big picture scholarly speculation. Naive models of ISO creation would suggest the maximal number of ISOs created by a system should be when it hosts the most free flying small planetesimals that are dynamically hot, that is, when the system is very young, within the first few Myrs of the host stars formation. But this is also the era when a system’s protoplanetary disk is optically thick near its midplane and planetesimals in the densest part of the disk remain cold and thermally unprocessed. PPDs are seen to clear within $5\mathrm{-}10\,\mathrm{Myrs}$ of host star formation though (Rieke et al., 2005; Ben et al., 2025, and references therein), and the first heavy thermal processing of KBOs estimated to occur in $1\mathrm{-}40\,\mathrm{Myrs}$ (Davidsson, 2021; Steckloff et al., 2021; Lisse et al., 2021, 2022b; Parhi & Prialnik, 2023). Planetary migration and Oort Cloud formation involving small body ejection processes can take many 100’s of Myrs to complete (Ref from Dearth papers). We can thus argue plausibly that PPDs can clear and small body thermal evolutionary processes can occur more rapidly than dynamical processes can cool down and mature in a system, and we should see a mix of volatile rich, hypervolatile depleted, and totally devolatilized ejected objects as our ISOs. This is precisely what we have seen so far in highly-devolatilized 1I/Oumuamua, hypervolatile-rich 2I/Borisov, and well-thermally processed 3I.
6 Conclusions and Future Work
[…]
Registration and co-adding of dozens of scattered light continuum images at 1.0 - 1.5 um produced a high S/N ratio image consistent with a stellar point source of no extension. The same images were used to create a lightcurve, for which variability $\lesssim 15\,\mathrm{\%}$ was seen from beginning to end of the observing period. The absolute brightness of 3I at 1.0 - 1.5 um is consistent with a nucleus of effective radius Rnuc $\lesssim$ 2.5 km, as reported by HST (Jewitt+ 2025) surrounded by a 100 times brighter coma. The SPHEREx spectral continuum measured over 2 weeks is highly consistent with the continuum reported by JWST during its single visit to the object on 06 Aug 2025 (Cordiner et al. 2025). The 1.5-4.0 um continuum structure shows a strong feature commensurate with water ice + refractory organics absorption. The overall shape of this reflectance continuum is very similar to that reported for the cliff-type KBOs by Pinilla-Alonso, N., Brunetto, R., De Prá, M.N. et al. 2025, “A JWST/DiSCo-TNOs portrait of the primordial Solar System through its trans-Neptunian objects”, Nat Astron 9, 230.
The August SPHERExobservations provided the first detection of strong $\mathrm{CO}_{2}$ emission from the comet into a giant extended surrounding coma, proving that $\mathrm{CO}_{2}$ sublimation from large icy coma grains is the main driver for the observed outgassing activity and explaining its extended morphology as the product of $\mathrm{CO}_{2}$ gas sublimation that was undetectable in the first reported UVIS characterizations of the object. The moderate observed water gas production rate can be explained as an object just coming in from the cold depths of interstellar space that was observed at rh = 3.2 au from the Sun outside the water ice line.
In many ways the observed activity behavior of 3I, including its preponderance of $\mathrm{CO}_{2}$ emission, lack of $\mathrm{CO}$ output, small size, and predominance of large chunks of water-ice rich icy material in a flux-dominant coma is reminiscent of the behavior of short period comet 103P, target of the NASA Deep Impact extended mission in 2010 and a “hyperactive comet” near the end of its outgassing lifetime (A’Hearn et al., 2011; Lisse et al., 2009). Such an object would have undergone hundreds to thousands of close perihelion passages by its natal birth star inside the system’s water ice line, as 103P has by our Sun. Whether this was due to 3I forming within a few au of its birth star or being scattered into this region after formation but before ejection is not clear. In terms of ice incorporation and thermal processing mechanisms operating on solar system comets (Lisse et al., 2021, 2022a), 3I is closer in evolution to barely- or non-active 1I/Oumuamua than primitive, ice rich 2I/Borisov.
Compared to the naïve expectation that solar systems shed the majority of their comets in their first few Myrs of existence when their small icy planetesimals are most abundant, most hypervolatile ice rich, and most dynamically hot, so that all ISOs should be like 2I/Borisov, this suggests that ISOs are thermally processed quicker than they dynamically relax. This hypothesis needs to be tested by population studies of many more ISOs than the 3 currently known, and hopefully many more will be studied by the SPHERExmission’s all-sky survey. Another inference that can be made about 3I’s home system is that it contains a companion body of mass MNeptune or larger within a few au of the host star; such a massive body is required to eject a body like 3I from a stellar system with host star of class M8V or larger.
This paper relies on data obtained by the SPHEREx Observatory (operated by Caltech and JPL on behalf of the National Aeronautics Space Administration under contract 80GSFC18C0011) that is hosted by IPAC as part of the IRSA archive (10.26131/IRSA652 SPHEREx Quick Release Spectral Images - QR2). We are indebted to the SPHEREx project for obtaining additional pointings at 3I, approximately 100 percent more, than their nominal survey would have obtained less than 3 months after the start of the survey by testing and implementing a new commanding protocol. Part of the research described in this paper was carried out at the Jet Propulsion Laboratory, California Institute of Technology, under a contract with the National Aeronautics and Space Administration (80NM0018D0004). The authors also acknowledge the Texas Advanced Computing Center (TACC)  [^10] http://www.tacc.utexas.edu at The University of Texas at Austin for providing computational resources that have contributed to the research results reported within this paper. {contribution}All listed authors contributed importantly to the creation and production of this manuscript.
Appendix A Appendix information